Systems of Equations
1. Solve the following problems for X and Y i these two system equations, and for X and Y and Z, in these three equations.
2. Make sure show your work step by step.
3. Fill the table:
Q# Equation 1: Equation 2 : Equation 3: X= : Y= : Z= :
________________________________________________________________________________
1. X+Y=7; X+2Y=11; none -- -- none
________________________________________________________________________________
2. 3X+7Y=24; 9X+5Y=24; none -- -- none
________________________________________________________________________________
3. 4X+Y=17; 2X+3Y=21; none -- -- none
________________________________________________________________________________
4. 18X+Y=42; 6X-3Y=-6; none -- -- none
________________________________________________________________________________
5. X+Y+2Z=5; X+Y=5; X+3Y+Z=11; -- -- --
________________________________________________________________________________
6. X+10Y+Z=52; 5X+Y+4Z=15; X+2Y-3Z=12; -- -- --
________________________________________________________________________________
7. 2X+2Y+4Z=10; 3X+3Y=15; 4X+12Y+4Z=44; -- -- -- ____________________________________________________________________________
4. What is the difference between system #5 and system #7?

Answer by Alan3354(69443)

7.
2X+2Y+4Z=10 --> x + y + 2z = 5
3X+3Y=15 --> x + y = 5
4X+12Y+4Z=44 --> x + 3y + z = 11
------------
Reducing often helps.
x + y + 2z = 5
x + y + 0z = 5
---------------- Subtract
2z = 0
z = 0
------
x + 3y = 11
x + y = 5
----------- Subtract
2y = 6
y = 3
------
x = 2
------
Do the others in a similar manner
----------------
Don't type in all those lines.